OpenStudy (anonymous):

The volume of a rectangular box with a square base remains constant at 1000cm^3 as the area of the base increases at a rate of 2 cm^2/sec. Find the rate at which the height of the box is decreasing when each side of the base is 19cm long. (do not round the answer)

OpenStudy (saifoo.khan):

V = base area * height. 1000 = A * h Take the derivative of both sides with respect to t using the chain rule and product rule: \[0 = \frac{dA}{dt}(h) +\frac{dh}{dt}(A)\]

OpenStudy (saifoo.khan):

We are given that dA/dt = 2 Side is 19 so base area = 19^2 = 361

OpenStudy (saifoo.khan):

one more ting. Vol = 1000 A*h = 1000 361*h = 1000 find h. Gives you height. insert the values in the product rule.

OpenStudy (anonymous):

so the height would be decreasing at a rate of 2.77cm^2/sec?

OpenStudy (anonymous):

well it looks like one of my choices are \[A. \frac{ 2000 }{ 6859 }cm/\sec\] \[B. \frac{ 2 }{ 361 }cm/\sec\] \[C.\frac{ 2000 }{ 130321 }cm/\sec\] \[D. \frac{ 1000 }{ 361 }cm/\sec\] So would it be D then???

OpenStudy (saifoo.khan):

No!! That just gives you the height. Insert the values in the product rule:

OpenStudy (anonymous):

so if I take \[361*\frac{ d }{ dt }\frac{ 1000 }{ 361 }+\frac{ 1000 }{ 361 }\frac{ d }{ dt }(361)=\frac{ d }{ dt }(1000)\]

OpenStudy (anonymous):

I still don't understand

OpenStudy (saifoo.khan):